Given the following data, answer the questions below the table: ---------------------------------------------------------------------------------------------------------------------- Data Unit sales        

  1. Given the following data, answer the questions below the table:

----------------------------------------------------------------------------------------------------------------------

Data

Unit sales

        30,000

units

Selling price per unit

$10

per unit

Variable expenses per unit

$5

per unit

Fixed expenses

$120,000

----------------------------------------------------------------------------------------------------------------------

a) What is the break-even in dollar sales?

b) What is the margin of safety percentage?

c) What is the degree of operating leverage? (Round your final answer to 2 decimal places.)

d) Using the degree of operating leverage, calculate the percentage change in net operating income if unit sales increase by 20%.
Answer #1

Contribution margin ratio = Contribution margin/Sales price

= (10-5)/10 = 50%

Breakeven in dollars = Fixed cost/Contribution margin ratio

= 120,000/50%

= 240,000

Margin of Safety = (Sales - Breakeven sales) /Sales

Sales = 30,000*10 = 300,000

= (300,000-240,000)/300,000

= 20%

Contribution margin = (10-5)*30,000 = 150,000

Net operating income = Contribution margin - Fixed cost

= 150,000 - 120,000 = 30,000

Degree of operating leverage = Contribution margin/Net operating income

= 150,000/30,000

= 5

Increase in net operating income = increase in sales* degree of operating leverage

= 20 * 5

= 100% increase